A bag of batteries in the ratio of 3 size D batteries to 7 size C batteries is combined with a bag having a ratio of 5 size D to 1 size C.
If the two bags together now contain an equal number of size D and C batteries, how many batteries are in the second bag?
My attempt:
Step 1: I can see that the first bag in the ratio of 3/7 has 18 size D batteries and 42 size C batteries, given that the total number in the bag is 60.
Step 2: I can see that the second bag is in the ratio of 5/1.
Step 3: Through trial and error, I can see that I need to get to a combined bag with a ratio of 8/8 for a total of 96. But I don't know how to do this mathematically.
This is often my problem. Can anybody help me set this up?
Thank you.

Answer by Theo(13342)

you are correct in that both bags will contain 96 batteries.
there will be 48 D and 48 C.

you are given that the first bag contains 60 batteries total and that the ratio of D to C is 3/7

let x = the common multiplier.

you get 3x + 7x = 60

combine like terms to get 10x = 60

solve for x to get x = 6

the common multiplier is 6.

3*6 + 7*6 = 18 + 42 = 60

the total number of batteries is 60.

the ratio of D to C is 18 / 42 = 3 / 7.

the batteries in the second bag are in the ratio of D / C = 5 / 1

let y be the common multiplier for the second bag.

you get 5y + 1y = z

z is the total for the second bag.

when you add the D batteries together and you add the C batteries together, you will get a ratio of 1/1.

you get:

(18 + 5y) / (42 + y) = 1/1 which becomes:

(18 + 5y) / (42 + y) = 1

multiply both sides of this equation by (42 + y) to get

18 + 5y = 42 + y

subtract y from both sides of this equation and subtract 18 from both sides of this equation to get:

5y - y = 42 - 18

simplify to get:

4y = 24

solve for y = get y = 6

when y = 6, the equation for your second bag of 5y + y = z becomes 30 + 6 = z which results in 36 = z

your second bag contains 36 batteries.
30 of them are D
6 of them are C.
the ratio of D/C = 30/6 = 5/1.

your combined bag contains 60 + 36 = 96 batteries.
your total D are 18 + 30 = 48
your total C are 42 + 6 = 48
the ratio of D to C is 1/1.